OpenStudy (anonymous):

what is the equation of a line perpendicular to 4x+8y=16

OpenStudy (anonymous):

A good way to start with these questions is to get it in the form. \[y = mx + c\]Can you do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

No problem, we just need to rearrange our equation. At the moment we have \[4x + 8y = 16\]We can subtract \(4x\) from both sides to give \[8y = 16 - 4x\]Now we can divide everything by 8, giving \[y = 2 - 0.5x\]With me so far?

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

Great! So you can see that the form \[y = mx + c\]is represented by our equation\[y = -0.5x + 2\]We have \(m = -0.5\) and \(c = 2\)

OpenStudy (anonymous):

Now another way of saying \(0.5\) is the same as \[1\over2\]right?

OpenStudy (anonymous):

So we can make our equation \[y = {-1x\over 2} + 2\]Now the gradient of this line is -1/2 which will look something like thisWith me so far?

OpenStudy (anonymous):

We haven't finished yet, but is there anything so far which doesn't make sense?

OpenStudy (anonymous):

I'm going to bed now (I'm in Australia) but the answer is \[y = 2x+c\]where c can be any real number. Because to find the perpendicular of a line you take the inverse reciprocal of the gradient. In our case\[{-1 \over 2} \rightarrow 2\]

OpenStudy (anonymous):

Hope that helped, please let me know if there is something you don't get.